OpenStudy (anonymous):

A disc is rotating as per the relation teta= at^2 + bt + c where a , b ,and c are constant t is the time and teta=angular displacement it is brought to rest in 10 rotations then b=?

OpenStudy (anonymous):

teta?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

\[\theta=angular displacement\]

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

A disc is rotating as per the relation teta= at^2 + bt + c where a , b ,and c are constant t is the time and teta=angular displacement it is brought to rest in 10 rotations then b=?

OpenStudy (anonymous):

What is the answer?

OpenStudy (anonymous):

\[\sqrt{-80apie}\]

OpenStudy (anonymous):

@Ishaan94 wat do u get

OpenStudy (anonymous):

At t=0, the disc is already rotating. So, theta = c. (theta for teta) After ten rotations theta=10*2pi. => 20pi = at^2 + bt + c. t is the time after 10 rotations. Is this all information you got? I am bit confused.

OpenStudy (anonymous):

Ohh theta is angular displacement. Sorry. I took it as the angular distance.

OpenStudy (anonymous):

I am still confused though.

OpenStudy (anonymous):

No Problem Bro just skip this and take a look at my New question plzz

OpenStudy (anonymous):

\[\theta(t)=at^2+bt+c\]\[\omega(t)=d \theta(t)/dt=2at+b\] when it stops: \[\omega(T)=0=2aT+b \rightarrow T=-b/2a\] Replacing in the angle equation: \[\theta(T)=a(-b/2a)^2+b(-b/2a)+c=c-b^2/4a\] \[\Delta \theta=\theta(T)-\theta(0)=c-b^2/4a-c=-b^2/4a=20\pi\] \[b=\sqrt{-80\pi a}\]